Financial data for Joel de Paris, Inc., for last year follow:

Joel de Paris, Inc.
Balance Sheet
Beginning
Balance
Ending
Balance
Assets
Cash $ 131,000 $ 128,000
Accounts receivable 333,000 485,000
Inventory 577,000 489,000
Plant and equipment, net 819,000 798,000
Investment in Buisson, S.A. 405,000 430,000
Land (undeveloped) 255,000 251,000
Total assets $ 2,520,000 $ 2,581,000
Liabilities and Stockholders' Equity
Accounts payable $ 378,000 $ 332,000
Long-term debt 975,000 975,000
Stockholders' equity 1,167,000 1,274,000
Total liabilities and stockholders' equity $ 2,520,000 $ 2,581,000


Joel de Paris, Inc.
Income Statement
Sales $ 3,948,000
Operating expenses 3,355,800
Net operating income 592,200
Interest and taxes:
Interest expense $ 111,000
Tax expense 191,000 302,000
Net income $ 290,200


The company paid dividends of $183,200 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%.

Required:

1. Compute the company's average operating assets for last year.

2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Round "Margin", "Turnover" and "ROI" to 2 decimal places.)

3. What was the company’s residual income last year?
Answer #1

Average operating assets

= [(2,520,000-255,000-405,000)+(2,581,000-251,000-430,000)/2

= 1,880,000

Margin = Net operating income / sales

= 592,200/3,948,000

= 15%

Turnover = Sales / Average operating assets

= 3,948,000/1,880,000

= 2.1

Roi = Margin* turnover

= 15*2. 1

= 31.50%

Residual income = Net operating income - Required return

= 592,200 - (1,880,000*15%)

= 310,200

Comment if you face any issues